OpenStudy (dmezzullo):

A car dealer offers you two deals on a car that costs $16,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $3,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 1% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time (in years). Add the amount of interest to the price of the car.) Which monthly payment amount is lower? Plea

OpenStudy (dmezzullo):

@mikala1

OpenStudy (dmezzullo):

@jazy

OpenStudy (anonymous):

im way ahead of u in school i did this monthis ago

OpenStudy (dmezzullo):

i know im in module 2 of segment 1

OpenStudy (dmezzullo):

@jhonyy9

OpenStudy (anonymous):

hold on i got the answer

OpenStudy (anonymous):

I miss algebra -_- Anyways, just solve for the monthly payment options. Payment Option 1) You can finance the car for 60 months with no interest if you make a $3,000 down payment. SO: 16,000 - 3,000 = 13,000 13,000 /60 = 216.666 Monthly payment: $216.67 Payment Option 2) You can finance the car for 72 months (6 years) with 1% simple annual interest and no down payment. SO: 16,000 divided into 6 years = 2666.67 Multiply by 1% interest... 0.1 * 2666.67 = 26.67 2666.67 + 26.67 = 2693.34 per year. Divide that by 12 months. 2693.34 /12 = 224.45 Monthly Payment: $224.45

OpenStudy (anonymous):

Which is cheaper?

OpenStudy (dmezzullo):

option a

OpenStudy (anonymous):

Yep.

OpenStudy (dmezzullo):

tyvm lol medal from everyone?

OpenStudy (anonymous):

i remeber this in 6th grade at the way end in like history idk y but i completely forgot this lol

OpenStudy (anonymous):

Yes A medal for everyone xD

OpenStudy (dmezzullo):

no everyone give me a medal lol

OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $16,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $3,000 down payment. Total Price of the Car = Car price + down payment to the financer Total Price of the Car = $ 16,000 + 3,000 ======================================… Total Price of the Car = $ 19,000 ◄ higher in price Ans ======================================… Payment Option 2: You can finance the car for 72 months (6 years) with 1% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time (in years). Add the amount of interest to the price of the car.) Interest payment = Principal * Rate * Time Interest payment = 16,000 * 0.01 * 6 Interest payment = 960 Total Price of the Car = Price of the Car + Interest Payment Total Price of the Car = $ 16,000 + $ 960 ======================================… Total Price of the Car = $ 16,960 ◄ The choice Plan Ans ======================================… Which monthly payment amount is lower? Please explain how you arrived at your answer and show all of your work for each option.

OpenStudy (anonymous):

smh...wow demezz that ain't right. -.- I took Algebra in 8th grade, last year...it kinda stuck with me. (:

OpenStudy (anonymous):

lol i am in 8th

OpenStudy (dmezzullo):

ive only gotten uno medal

OpenStudy (anonymous):

lol i got no medals it is fine xP

OpenStudy (anonymous):

Wow, and I'm supposedly advanced, lol! You're really smart then @Converse_Luv :)

OpenStudy (anonymous):

can u guys help me please ?

OpenStudy (anonymous):

lol thanks x) yay medal xD

OpenStudy (anonymous):

jazy and anyone

OpenStudy (anonymous):

sure link?

OpenStudy (anonymous):

kk hold on :)

OpenStudy (anonymous):

omg i have to take an english exam all o er again